OpenStudy (anonymous):

How do you convert 1.42 g/cm^2 to mg/mm^2 Step by Step?

OpenStudy (mkmkasim):

the answer is 14.2 mg/mm^2

OpenStudy (mkmkasim):

first you multiply by 1000 to convert to mg.

OpenStudy (mkmkasim):

then you dived by (100^2) to convert to mm^2

OpenStudy (anonymous):

Thanks so much! I knew the answer just not how to get it. ! (:

OpenStudy (mkmkasim):

No problem!

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \begin{array}{llll} 1cm&10mm\\ 1cm^2&10^2mm^2 \end{array}}}\\ \quad \\ \cfrac{1.42\cancel{g}}{\cancel{cm^2}}\cdot \cfrac{\cancel{cm^2}}{10^2mm^2}\cdot \cfrac{1000mg}{\cancel{g}}\)